# Segmented Operations using Matrix Multiplications

### v1

@@ Abstract
Specialized computational units that perform small matrix multiplications as primitive operations are typically present in modern accelerators. However, these units are often underutilized for many fundamental operations besides dense matrix multiplications. The analysis of algorithms for such architectures is currently stagnated due to the lack of a rigorous theoretical model of computation that captures their characteristics. In this work, we propose MMV-RAM, a computational model tailored to matrix multiplication accelerators. MMV-RAM judiciously extends the Vector-RAM model with an additional processing unit that multiplies two matrices of sizes $n\times s$ and $s\times s$ in a single parallel step, where $s$ is a model parameter. We provide a detailed theoretical analysis of the model, and carefully balance the computational power between the matrix and vector units, guided by the circuit complexity lower bound that parity is not in AC[0].
In MMV-RAM, we study algorithms for segmented scan and sum, two fundamental parallel primitives. We propose a segmented scan algorithm that uses matrix multiplications to perform speculative block-scan computations, which runs in $\mathcal{O}(\log_{s}(n))$ steps. In contrast, we show that any algorithm that uses only the vector unit of MMV-RAM requires $\Omega\left(\frac{\log_{2}(n)}{\log_{2}\log_{2}(n)}\right)$ steps. We further apply these techniques to obtain similar theoretical speedups for element-wise vector multiplication and matrix multiplication. Beyond the worst-case complexity analysis, we propose algorithms for segmented operations that could lead to highly efficient and pragmatic implementations. For example, we observe that segmented sum is a combination of three elementary parallel primitives: scan, compress, and vector differentiation. As a case study, we implement and experimentally evaluate the proposed algorithms using the Ascend AI 910B accelerator, which contains both matrix multiplication and vector cores, demonstrating significant speedups against a vector-only baseline.
@@ 1 Introduction
In recent years, the unexpectedly rapid evolution and development of Artificial Intelligence (AI) and, in particular, deep learning have strongly influenced the computing industry. Deep learning workloads, both training and inference, demand more and more performance every year [31]. The high demand for accelerated AI computing has resulted in manufacturing domain-specific accelerators that contain, most notably, specialized Matrix Multiplication Unit (MMU). Examples of such accelerators and MMU that have been massively produced include NVIDIA Tensor Cores [43], Google TPU Matrix Multiplication unit [30], and Huawei Ascend’s Cube unit [35], to name a few. These MMU have helped to increase the peak arithmetic performance for regular and dense computations.
@@
Although MMU have proven highly effective for regular and dense computations, such as convolutions, the attention mechanism in transformer architectures [56], and fully connected layers in deep learning, their applicability to irregular and sparse computations remains uncertain. In particular, sparsity in deep learning is a well-known characteristic that has not yet been successfully exploited in practice [25, 18, 13]. Therefore, there is a debate in the computing industry on whether specialized computing units for sparse computations, usually referred to as “Sparse Tensor Cores”, should be designed [58, 42, 24], or if irregular and sparse computations should be handled using the already existing MMUs along with algorithmic innovations [14, 15, 63, 44, 33]. In this work, we follow the latter approach of developing algorithmic innovations using MMU as initiated in [15] and revisit segmented operations (scan and sum) in this new setting. The motivation why we focus on segmented operations is that, first, they are fundamental computational paradigms in parallel computing with various applications [6, 49, 12]; second, the most prominent application of segmented sums is irregular sparse matrix-vector multiplication [9, 37] which is of great importance to efficiently address irregular and sparse computations in the near future.
@@
The analysis of parallel algorithms, especially for prefix sum or scan problems, is well-established in parallel models of computation such as the Parallel RAM (PRAM) [50, 28, 20], Bulk Synchronous Parallel (BSP) [55, 39, 40, 54], LogP [16], Vector-RAM [46, 7], the TCU Model (the first that tries to model matrix multiplication accelerators) [14, 15], and circuit models [20, 57, 59, 34]. As we target AI Accelerators, containing both matrix and vector units, the two most relevant models are arguably TCU and Vector-RAM, but none of them fully embrace the accelerator heterogeneity, as we discuss next.
@@
The TCU model is not suitable for our purposes because it omits any form of vector processing unit. This omission is both technically non-trivial to address (as we discuss extensively in this work), and essential for accurately modeling modern architectures. Indeed, Vector Unit (VCU) are typically present near the matrix multiplication units in AI accelerators, since element-wise activation operations are ubiquitous in AI workloads. For example, in convolutional neural networks, convolutions are followed by the ReLu activation function [32], and, in the transformer architecture, the linear layers of FFN are followed by the specialized GELU activation function [56].
@@
The Vector-RAM model, instead, also presents some limitations in our context. First, it does not entail an explicit representation of matrix multiplication units, making it inadequate for reasoning about algorithms that depend heavily on such operations. Moreover, its vector instruction set is overly powerful: they can perform operations like prefix sums, or even matrix multiplication, in a single step [7]. With such a powerful vector unit, an explicit matrix multiplication unit becomes obsolete.
@@ 1.1 Contributions
Our key technical contribution in the design of the MMV-RAM model is the separation and balancing of computational power between the matrix and vector units, ensuring that they complement rather than subsume each other. This balance is grounded in a fundamental result from circuit complexity: the parity function cannot be computed by constant-depth, polynomial-size (denoted by AC[0]) circuits [21, 62, 22]. Therefore, as long as the vector unit of MMV-RAM is restricted to AC[0], then it cannot efficiently simulate the matrix unit, thus maintaining a clear computational distinction between the two units.
MMV-RAM adopts a standard work/depth cost framework, enabling an intuitive analysis of algorithms and overcoming the restrictive external-memory model view of the TCU model [15, Section $5$]. For algorithm engineers, it is straightforward to use MMV-RAM for algorithmic analysis: the step complexity is given by the number of matrix and vector operations, while the total work is the sum of the costs of these operations.
Given the definition of the MMV-RAM model, we next design and analyze algorithms, starting with segmented operations. Algorithm 1 is an MMV-RAM segmented scan algorithm that executes (speculative) unsegmented scan operations using the MMU, and then efficiently corrects the mispeculated scans with the vector unit. The following theorem, which is proved in Section 4.3, summarizes its analysis.
Theorem 1.1 (Informal version of Theorem 4.4).
Given a value vector $\bm{\mathrm{x}}$ with bounded integer elements $|\bm{\mathrm{x}}(i)|\leq M$, and a boolean vector $\bm{\mathrm{f}}$, both of size $n$, Algorithm 1 computes the segmented scan $\bm{\mathrm{z}}$ of $\bm{\mathrm{x}}$ over $\bm{\mathrm{f}}$. Assuming that in MMV-RAM a single matrix multiplication between two $n\times s$ and $s\times s$ matrices requires $\mathcal{O}(ns^{2})$ operations (work), the algorithm requires:
  • $\mathcal{O}(\log_{s}(n))$ steps,
  • $\mathcal{O}\left(nB(s+\frac{B}{s})\right)$ total work,
where $B$ is the number of bits-per-element. On the contrary, any algorithm that uses only the vector unit and executes work that scales polynomially in $n$ requires $\Omega\left(\frac{\log(n)}{\log(\log(n))}\right)$ steps.
In summary, our main contributions are the following:
@@
    The MMV-RAM model of computation is proposed as an extension of the Vector-RAM model that contains both matrix multiplication and vector computing units, targeting matrix multiplication / AI accelerators. We provide a detailed analysis of the model, carefully balancing the power of the two units by exploiting the classic parity lower bound from circuit complexity.
@@
    In the MMV-RAM model, we design an $\mathcal{O}(\log_{s}(n))$-depth segmented scan algorithm (Algorithm 1) and analyze two fundamental applications, namely, element-wise vector multiplication and matrix multiplication. Our analysis indicates that by exploiting the MMU, we can achieve a provable theoretical speedup, compared to any MMV-RAM algorithm that uses only vector instructions.
@@
    Furthermore, we design multi-core speculative algorithms for segmented scan/sum, (Algorithms 1 and 2), as straightforward compositions of a few fundamental parallel primitives, applying speculation over the full input range. These algorithms are implemented on the Ascend 910B AI accelerator, which features matrix multiplication and vector processing cores. By leveraging speculative execution and effectively utilizing the matrix cores, we achieve substantial performance gains compared to vector-only baselines.
@@ 1.2 Notation
Matrices and vectors are denoted by capital and small letters, respectively, in bold font. We denote by $\bm{\mathrm{U}}_{s}$ the upper-triangular all-ones square matrix of size $s$ (including the ones on the diagonal). All vectors are considered column vectors, and $\bm{\mathrm{A}}^{\top}$ and $\bm{\mathrm{x}}^{\top}$ denote the transpose of matrix $\bm{\mathrm{A}}$ and vector $\bm{\mathrm{x}},$ respectively. Vectors and matrices are zero-index based. For a vector $\bm{\mathrm{x}}$, $\bm{\mathrm{x}}(i)$ is the $(i+1)$-th entry, and for integers $i<j$, we denote by $\bm{\mathrm{x}}(i:j)$ the subvector $(\bm{\mathrm{x}}(i),\bm{\mathrm{x}}(i+1),\dots,\bm{\mathrm{x}}(j-1))$. $\bm{\mathrm{x}}(i:j:s)$ is the $s$-strided subvector with offset $i$: $(\bm{\mathrm{x}}(i),\bm{\mathrm{x}}(i+s),\bm{\mathrm{x}}(i+2s)\dots,\bm{\mathrm{x}}(j-1))$. If $j$ is omitted, that is, when we write $\bm{\mathrm{x}}(i::s)$, then it defaults to the length of $\bm{\mathrm{x}}$. For a vector $\bm{\mathrm{x}}$ of length $n$, the operation $\bm{\mathrm{z}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{x}},\bm{\mathrm{A}})$ denotes that $\bm{\mathrm{x}}$ is viewed as a matrix $\bm{\mathrm{X}}$ of size $\lceil n/s\rceil\times s$, in a row-major order (zero-padded if necessary), and the result $\bm{\mathrm{z}}$ is a vector containing the elements of the matrix product $\bm{\mathrm{X}}\bm{\mathrm{A}}$, in row-major order as well. “Hat” accents on vectors, e.g. $\widehat{\bm{\mathrm{x}}}$, denote the scanned version of the vectors. The “bar” accent typically denotes that all consecutive blocks of length $s$ are prefix-summed within each block, e.g., $\overline{\bm{\mathrm{x}}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{x}},\bm{\mathrm{U}}_{s})$. All logarithms are base two, unless explicitly specified. For a matrix $\bm{\mathrm{A}}$, $\|\bm{\mathrm{A}}\|_{\max}$ is the maximum absolute value over all matrix elements.
@@ 1.3 Outline
In Section 2, we recall some basic definitions and background. In Section 3 we describe the MMV-RAM model. The main algorithms for segmented operations and their applications are described in Sections 4, 5, and 6. In Section 7, we provide an experimental evaluation of the proposed algorithms as a case study of the MMV-RAM model on Ascend accelerators. Section 8 outlines potential extensions and/or variants of the MMV-RAM model. We finally conclude in Section 9.
@@ 2 Background and definitions
In this section, we define the segmented scan and segmented sum tasks. Furthermore, we discuss a block-recursive approach to computing segmented scans. In the (unsegmented) scan task, we are given an array $\bm{\mathrm{x}}$ of length $n$ and the task is to return an array $\bm{\mathrm{z}}$ of length $n$ such that $\bm{\mathrm{z}}(i)=\sum_{j\leq i}\bm{\mathrm{x}}(j)$ for each $i$. The segmented scan task is a generalization of the scan task, where the input is divided into multiple segments, and the scan operation is performed independently within each segment:
@@
$\bm{\mathrm{z}}(i)=\begin{cases}\bm{\mathrm{x}}(i),&\text{if }i=0\text{ or }\bm{\mathrm{f}}(i)=1\text{ (segment start)}\\ \bm{\mathrm{z}}(i-1)+\bm{\mathrm{x}}(i),&\text{otherwise}.\end{cases}$ (1)
Segmented sum is defined accordingly in 2.2. It is implicitly assumed that $\bm{\mathrm{f}}(0)=1$.
Problem 2.2 (Segmented Sum).
Given the tuple $(\bm{\mathrm{x}},\bm{\mathrm{f}})$ as in 2.1, let $N$ be the number of segments, $\{p_{0}=0,p_{1},p_{2},\ldots,p_{N-1}\}$ be their starting indices, and $p_{N}=n$. The segmented sum (or reduction) of $\bm{\mathrm{x}}$ over $\bm{\mathrm{f}}$ is the array $\bm{\mathrm{z}}=\left(\bm{\mathrm{z}}(0),\bm{\mathrm{z}}(1),\ldots,\bm{\mathrm{z}}(N-1)\right)$, where $\bm{\mathrm{z}}(i)=\bm{\mathrm{x}}(p_{i})+\ldots+\bm{\mathrm{x}}(p_{i+1}-1)$.
An example of a segmented scan is shown in Equation  (2):
$\bm{\mathrm{x}}=\begin{pmatrix}2,2,3,3,1,3,1,2\end{pmatrix},\quad\bm{\mathrm{f}}=\begin{pmatrix}1,0,1,0,0,1,0,0\end{pmatrix},\quad\bm{\mathrm{z}}=\begin{pmatrix}2,4,3,6,7,3,4,6\end{pmatrix}.$ (2)
The corresponding segmented sum of this example would be $\bm{\mathrm{z}}=\left(4,7,6\right)$.
In the literature, scan problems typically assume a binary associative operator, i.e., a semi-group structure (see, e.g., [6, 20]). The algebraic properties of the binary scan operator that are required from our algorithms for segmented operations are more restrictive due to speculative execution. In particular, we require that the set of input elements with the corresponding binary operation form a group, since we require an identity element and inverses to revert speculative operations.
We will also use the compaction (also known as compress) operation. Given $\bm{\mathrm{x}}$ and $\bm{\mathrm{f}}$, compress returns all entries of $\bm{\mathrm{x}}$ whose corresponding entry on $\bm{\mathrm{f}}$ is set to true/one. For the example of Equation 2, $\text{COMPRESS}(\bm{\mathrm{x}},\bm{\mathrm{f}})$ returns $(2,3,3)$.
2.1 Model definitions: Vector-RAM, TCU, and circuits
Here, we provide some background on the models of computation that we will discuss throughout the paper.
Circuits
A circuit is a Directed Acyclic Graph (DAG) $G=(V,E)$. Nodes with zero in-degree are called inputs, and nodes with zero out-degree are called outputs. $|E|$ is the size of the circuit, whereas its depth is the longest path from an input to an output. The fan-in is the largest in-degree, and the fan-out is the largest out-degree. The vertices are also called gates, and they perform mathematical operations on their predecessors. In the scope of this work, the following types are considered:

### v2

@@ Abstract
Specialized computational units that perform small matrix multiplications as primitive operations are typically present in modern AI accelerators. However, these Matrix Multiplication Units are often underutilized for many fundamental deep learning operations besides dense matrix multiplications. Coincidentally, the lack of a rigorous theoretical model of computation for such architectures obstructs algorithmic design. In this work, we propose MMV-RAM, a computational model which judiciously extends the Vector-RAM model with an additional MMU. We provide a detailed theoretical analysis and carefully balance the computational power between the matrix and vector units, guided by the circuit complexity lower bound that parity is not in AC[0]. Given MMV-RAM, we proceed to algorithm design, starting with two fundamental parallel operations: segmented scan and sum. By expressing them as compositions of elementary parallel primitives (e.g., seg. sum reduces to: scan, compress, and vector differentiation), we can exploit MMUs to perform speculative blocked computations, ultimately leading to provable theoretical speed-ups against vector-only approaches. These results extend to other ubiquitous AI kernels, including dense matrix product, and sparse matrix-vector product. As a case study, we implemented the proposed algorithms on the Ascend 910B AI accelerator, which contains matrix and vector cores. We evaluate these implementations on synthetic and real-world datasets from various applications, including Large Language Models.
@@ 1 Introduction
In recent years, the unexpectedly rapid evolution and development of Artificial Intelligence (AI) and, in particular, deep learning have strongly influenced the computing industry. Deep learning workloads, both training and inference, demand more and more performance every year [34]. The high demand for accelerated AI computing has resulted in manufacturing domain-specific accelerators (often referred as Neural Processing Units (NPUs)) that contain, most notably, specialized MMUs. Examples of such accelerators and MMUs that have been massively produced include NVIDIA Tensor Cores [47], Google TPU Matrix Multiplication unit [33], and Huawei Ascend’s Cube unit [38], to name a few.
@@
Although MMUs have proven highly effective to increase the peak arithmetic performance for regular and dense computations, such as convolutions, the attention mechanism in transformer architectures [60], and fully connected layers in deep learning, their applicability to irregular and sparse computations remains uncertain. In particular, sparsity in deep learning is a well-known characteristic that has not yet been successfully exploited in practice [27, 18, 13]. Therefore, there is a debate in the computing industry on whether specialized computing units for sparse computations, usually referred to as “Sparse Tensor Cores”, should be designed [62, 46, 26], or if irregular and sparse computations should be handled using the already existing MMUs along with algorithmic innovations [14, 15, 67, 48, 36].
In this work, we follow the latter approach, exploiting MMUs as initiated in [15] and revisit segmented operations (scan and sum) in this new setting. Our motivation to focus on segmented operations is two-fold. First, they are fundamental computational primitives in parallel computing with various applications [6, 53, 12]; second, the most prominent application of segmented sums is irregular sparse matrix-vector multiplication [9, 40] which is of great importance to efficiently address irregular and sparse computations in the near future.
@@
The analysis of parallel algorithms, especially for prefix sum or scan problems, is well-established in parallel models of computation such as the Parallel RAM (PRAM) [54, 31, 21], Bulk Synchronous Parallel (BSP) [59, 43, 44, 58], LogP [16], Vector-RAM [50, 7], the TCU Model (the first that tries to model matrix multiplication accelerators) [14, 15], and circuit models [21, 61, 63, 37]. As we target AI accelerators, containing both matrix and vector units, the two most relevant models are arguably TCU and Vector-RAM, but none of them fully embraces the accelerator heterogeneity, as we discuss next.
@@
The TCU model is not suitable for our purposes because it omits any form of vector processing unit. This omission is both technically non-trivial to address (as we discuss extensively in this work), and essential for modeling modern architectures. Indeed, Vector Units are typically present near the matrix multiplication units in AI accelerators, since element-wise activation operations are ubiquitous in AI workloads. For example, in convolutional neural networks, convolutions are followed by the ReLu activation function [35], and, in the transformer architecture, the linear layers of FFN are followed by the specialized GELU activation function [60].
@@
The Vector-RAM model, instead, also presents some limitations in our context. First, it does not entail an explicit representation of MMUs, making it unbefitting for reasoning about algorithms that depend heavily on such operations. Moreover, its vector instruction set is overly powerful: VCUs can perform operations like prefix sums, or even matrix multiplication, in a single step. With such a powerful vector unit, an explicit matrix multiplication unit becomes obsolete.
@@ 1.1 Contributions
Our key technical contribution in the design of MMV-RAM is the separation and balancing of computational power between the matrix and vector units, ensuring that they complement rather than subsume each other. This balance is grounded in a fundamental result from circuit complexity: the parity function cannot be computed by constant-depth, polynomial-size (denoted by AC[0]) circuits [22, 66, 24]. Therefore, as long as the vector unit of MMV-RAM is restricted to AC[0], then it cannot efficiently simulate the matrix unit, thus maintaining a clear computational distinction between the two units.
MMV-RAM adopts a standard work/depth cost framework, enabling an intuitive analysis of algorithms. For algorithm engineers, it is straightforward to use MMV-RAM for algorithmic analysis: the number of steps (or “time complexity”) of an algorithm is the number of matrix and vector operations, while the total work is the sum of their corresponding costs.
Given the definition of the MMV-RAM model, we next design and analyze algorithms, starting with segmented operations. We first observe that these operations can be written as compositions of elementary parallel primitives, which lead to our main Algorithms 2 and 3. Our analysis indicates that by exploiting the MMU of MMV-RAM we can obtain provable theoretical speed-ups against vector-only implementations. Our main resuls are summarized in the following informal theorem.
Theorem 1.1 (Informal, see Thms. 4.2 and 4.3).
In MMV-RAM, there exist algorithms for segmented sum and scan of length $n$, which require:
$\mathcal{O}(\log_{s}(n))\text{ steps, and }\mathcal{O}\left(nB(s+\tfrac{B}{s})\right)\text{ work},$
where $B$ is the number of bits-per-element. In contrast, any algorithm that uses only vector operations and executes work that scales polynomially in $n$ requires $\Omega\left(\frac{\log(n)}{\log(\log(n))}\right)$ steps.
As direct applications, we extend these results for higher level algorithms, using segmented operations as subroutines. The first example is the element-wise multiplication of two vectors with integer elements. It is well-known that integer multiplication is not in AC[0], since parity reduces to it. As such, this operation cannot be directly implemented on the VCU as a primitive instruction. But it can be efficiently implemented by exploiting both the MMU and VCU using segmented operations. Ultimately, we also discuss algorithms for the product of two matrices, and well as the product between a sparse matrix and a dense vector (SpMV). For all these applications we report theoretical speed-ups similar to those of Theorem 1.1.
Our contributions are summarized as follows:
@@
    The MMV-RAM model of computation is proposed as an extension of the Vector-RAM model, that contains both matrix and vector processing units, targeting matrix multiplication / AI accelerators. We analyze the model in detail, balancing the power of the two units by exploiting lower bounds from circuit complexity.
@@
    In MMV-RAM, we design $\mathcal{O}(\log_{s}(n))$-depth algorithms for segmented operations (Algs. 2, 3). We prove that by exploiting the MMU and using speculative operations, we can achieve theoretical speed-ups compared to any MMV-RAM algorithm that uses only vector instructions. We extend these results to other applications, notably, matrix product and sparse matrix-dense vector multiplication (SpMV).
@@
    Furthermore, we implement and experimentally evaluate our algorithms on the Ascend 910B AI accelerator, which features both matrix multiplication and vector processing cores. By leveraging speculative execution and effectively utilizing the matrix cores, we report substantial performance gains for segmented scan compared to vector-only baselines.
@@ 1.2 Notation
Matrices and vectors are denoted by capital and small letters, respectively, in bold font. All logarithms are base two, unless explicitly specified. For a matrix $\bm{\mathrm{A}}$, $\|\bm{\mathrm{A}}\|_{\max}$ is the maximum absolute value over all matrix elements. We denote by $\bm{\mathrm{U}}_{s}$ the upper-triangular all-ones square matrix of size $s$ (including the ones on the diagonal). All vectors are considered column vectors, and $\bm{\mathrm{A}}^{\top}$ and $\bm{\mathrm{x}}^{\top}$ denote the transpose of matrix $\bm{\mathrm{A}}$ and vector $\bm{\mathrm{x}},$ respectively. Vectors and matrices are zero-index based. For a vector $\bm{\mathrm{x}}$, $\bm{\mathrm{x}}(i)$ is the $(i+1)$-th entry, and for integers $i<j$, we denote by $\bm{\mathrm{x}}(i:j)$ the subvector $(\bm{\mathrm{x}}(i),\bm{\mathrm{x}}(i+1),\dots,\bm{\mathrm{x}}(j-1))$. $\bm{\mathrm{x}}(i:j:s)$ is the $s$-strided subvector with offset $i$: $(\bm{\mathrm{x}}(i),\bm{\mathrm{x}}(i+s),\bm{\mathrm{x}}(i+2s)\dots,\bm{\mathrm{x}}(j-1))$. If $j$ is omitted, that is, when we write $\bm{\mathrm{x}}(i::s)$, then it defaults to the length of $\bm{\mathrm{x}}$. For a vector $\bm{\mathrm{x}}$ of length $n$, the operation $\bm{\mathrm{z}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{x}},\bm{\mathrm{A}})$ denotes that $\bm{\mathrm{x}}$ is viewed as a matrix $\bm{\mathrm{X}}$ of size $\lceil n/s\rceil\times s$, in a row-major order (zero-padded if necessary), and the result $\bm{\mathrm{z}}$ is a vector containing the elements of the matrix product $\bm{\mathrm{X}}\bm{\mathrm{A}}$, in row-major order as well. “Hat” accents on vectors, e.g. $\widehat{\bm{\mathrm{x}}}$, denote the scanned version of the vectors. The “bar” accent typically denotes that all consecutive blocks of length $s$ are prefix-summed within each block, e.g., $\overline{\bm{\mathrm{x}}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{x}},\bm{\mathrm{U}}_{s})$.
@@ 1.3 Outline
In Section˜2, we recall some basic definitions and background. Section˜3 contains the analysis of the MMV-RAM model. The main algorithms for segmented operations and their applications are described in Section˜4. In Section˜5, we provide an experimental evaluation of the proposed algorithms, before concluding in Section˜6.
@@ 2 Background and definitions
In this section, we define the segmented scan and segmented sum tasks. In the (unsegmented) scan task, we are given an array $\bm{\mathrm{x}}$ of length $n$ and the task is to return an array $\bm{\mathrm{z}}$ of length $n$ such that $\bm{\mathrm{z}}(i)=\sum_{j\leq i}\bm{\mathrm{x}}(j)$ for each $i$. The segmented scan task is a generalization of the scan task, defined as follows:
@@
$\bm{\mathrm{z}}(i)=\begin{cases}\bm{\mathrm{x}}(i),\qquad\qquad\text{if }i=0\text{ or }\bm{\mathrm{f}}(i)=1,\\ \bm{\mathrm{z}}(i-1)+\bm{\mathrm{x}}(i),\text{ otherwise}.\end{cases}$ (1)
Segmented sum is the collection of the last elements of each segment of the segmented scanned array. As an example, assume $\bm{\mathrm{x}}=(2,2,3,3,1,3,1,2)$ and $\bm{\mathrm{f}}=(1,0,1,0,0,1,0,0)$. The segmented scan and sum are $(2,4,3,6,7,3,4,6)$ and $(4,7,6)$, respectively.
In the literature, scan problems typically assume a binary associative operator, i.e., a semi-group structure (see, e.g., [6, 21]). Here, our algorithms require that the set of elements and the scan operator must form a group.
Last, we define compress that returns all entries of $\bm{\mathrm{x}}$ whose corresponding entry on $\bm{\mathrm{f}}$ is set to true/one. For the example above, $\text{COMPRESS}(\bm{\mathrm{x}},\bm{\mathrm{f}})$ returns $(2,3,3)$.
2.1 Circuits
A circuit is a Directed Acyclic Graph (DAG) $G=(V,E)$. Nodes with zero in-degree are called inputs, and nodes with zero out-degree are called outputs. $|E|$ is the size of the circuit, whereas its depth is the longest path from an input to an output. The fan-in is the largest in-degree, and the fan-out is the largest out-degree. The vertices are also called gates, and they perform mathematical operations on their predecessors.